OpenStudy (anonymous):

x+21 10 ----- = ------ 10 x find x?

sam (.sam.):

\[\begin{array}{l} \text{Multiply both sides by }x \\ \frac{1}{10} x (x+21)=10 \\ \text{Expand out terms of the left hand side:} \\ \frac{x^2}{10}+\frac{21 x}{10}=10 \\ \text{Divide both sides by }\frac{1}{10} \\ x^2+21 x=100 \\\end{array}\] \[\text{Can you continue?}\]

OpenStudy (anonymous):

I got there by u wasn't sure how to finish

sam (.sam.):

you need to make it like a quadratic equation then factor, \[x^2+21x=100\] \[x^2+21x-100=0\] \[(x+25)(x-4)=0\] Set each bracket =0 and solve for x

OpenStudy (anonymous):

oh, ok. so x=4 because length cant be negative?

sam (.sam.):

yes

OpenStudy (anonymous):

thanks

sam (.sam.):

welcome